OpenStudy (anonymous):

Find the value of k so that a line passing through (k, k – 3) and (–2, 4) has a slope of 1/4. A. –3 B. 2 C. 5 D. 10

OpenStudy (mathstudent55):

Do you know how to find the slope of a line given two points on the line?

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

use this eq \[(y2-y1)=m(x2-x1)\] here y2=4 y1=k-3 x2=-2 x1=k and m=slope=1/4 put values and ans=10 option C.

OpenStudy (mathstudent55):

For points \((x_1, y_1)\) and \((x_2, y_2) \), the slope of the line that passes through the points is \(m = \dfrac{y_2 - y_1}{x _2 - x_1} \)

OpenStudy (mathstudent55):

If you look at the slope formula above, finding the slope is simple. First you subtract the y-coordinates of the 2 points. Then you subtract the x-coordinates of the two points. Then you divide the difference in y by the difference in x. You are given two points and a slope. Let's use the slope formula with your info.

OpenStudy (mathstudent55):

\(m = \dfrac{y_2 - y_1}{x_2 - x_1} \) \(\dfrac{1}{4} = \dfrac{(k - 3) - (4)}{(k) - (-2)} \) Do you understand the line above?

OpenStudy (anonymous):

Uhhh....not really. I'm sorry.

OpenStudy (mathstudent55):

I used the formula for the slope. I wrote the slope formula just above the last line. The left side is m, slope. Your slope is given as 1/4, so we have 1/4 on the left side. Ok?

OpenStudy (anonymous):

Okay

OpenStudy (mathstudent55):

On the right side, we have a fraction. In the numerator, we a subtracting the y-values of the two points. In the denominator, we a subtracting the x-values of the two points. That's how we find a slope.

OpenStudy (anonymous):

Thank you!

OpenStudy (mathstudent55):

Now you need to simplify the equation we have and solve it for k.

OpenStudy (mathstudent55):

\(\dfrac{1}{4} = \dfrac{(k - 3) - (4)}{(k) - (-2)}\) Step 1. Get rid of parentheses: \(\dfrac{1}{4} = \dfrac{k - 3 - 4}{k +2}\) Step 2. Combine like terms: \(\dfrac{1}{4} = \dfrac{k - 7}{k +2}\) Step 3.Cross multiply (1)(k + 2) = 4(k - 7) k + 2 = 4k - 28 -3k = -30 k = 10

OpenStudy (mathstudent55):

Now let's check For k = 10, the point is (k, k - 3) = (10, 10 - 3) = (10, 7) Does the line through (10, 7) and (-2, 4) have a slope of 1/4? \(m = \dfrac{7 - 4}{10 - (-2)} = \dfrac{3}{10 + 2} = \dfrac{3}{12} = \dfrac{1}{4} \) The slope is 1/4, so our answer is correct.

OpenStudy (anonymous):

Thank you so much!!!!

OpenStudy (mathstudent55):

You're welcome.